# Bayesian WeakS-to-Strong from Text Classification to Generation

### v1

@@ 5.1 Datasets and Models
For slot filling, Llama-2-7b [27]666https://huggingface.co/meta-llama/Llama-2-7b-hf was used as the strong model which yielded better performance than Qwen-7B in this task. The same set of weak models was used as the classification task. As before, both weak and strong models are finetuned with all model parameters.
5.2 Metrics
@@
The classification task is evaluated by accuracy and the SLU-F1 [5] is used to evaluate the performance of slot filling, which combines both word-level and character-level F1 scores to give partial credit to non-exact match predictions. Performance gap recovered (PGR) [7] is used to measure the fraction of the performance gap recovered with weak supervision, which is defined as follows:
@@ 6.1 Text Classification
| Single Weak-to-Strong | GPT2-Large | 0.816 | 0.547 | 0.832 | 0.635 |
| OPT-1.3B | 0.811 | 0.563 | 0.828 | 0.648 | |
| Pythia-1.4B | 0.774 | 0.418 | 0.800 | 0.540 | |
| Naive Multi-Weak | 0.818 | 0.593 | 0.844 | 0.726 | |
| Bayesian Mutli-Weak | 0.822 | 0.614 | 0.855 | 0.781 | |
The proposed Bayesian WeakS-to-Strong approach was first evaluated on a classification task. Table 1 shows performance of the strong model and the weak models trained using ground-truth labels, with the former being the ceiling of the Weak(S)-to-Strong approaches. The strong model has about 7 times the number of parameters as the weak models, which also leads to about 28% relative improvement in the classification accuracy. Results of Weak(S)-to-Strong approaches are shown in Table 2. in Eqn. (1) and Eqn. (5) were set to 0 if auxiliary loss was not used. When a single weak teacher was involved to train the strong model, about 50% of the strong performance was recovered (shown by PGR). The addition of auxiliary loss boosted the performance for all three weak models where the strong model tended to gradually rely on its own prediction as the training progresses.
The Naive multi-weak approach and Bayesian multi-weak using EDL were applied to the classification task for WeakS-to-Strong. It can be seen that the auxiliary loss was effective for both. The Naive multi-weak method raised the average PGR to 0.726 and Bayesian multi-weak further boosted PGR to 0.781. Furthermore, a p-test between the results of single Weak-to-Strong on GPT2-Large (with an accuracy of 0.832) and Naive multi-weak (with an accuracy of 0.844) gave a p-value larger than 0.2, while the GPT2-Large-to-Strong and Bayesian multi-weak approach (with an accuracy of 0.855) resulted in a p-value less than 0.05. The results showed the effectiveness of Bayesian estimation that took the confidence of each weak label into account.
@@ 6.2 Text Generation
| Strong Model (ceiling) | Llama-2-7B | 6.7B | 0.781 |
@@
| Single Weak-to-Strong | GPT2-Large | 0.673 | 0.108 | 0.666 | 0.047 |
| OPT-1.3B | 0.634 | -0.265 | 0.569 | -0.831 | |
| Pythia-1.4B | 0.695 | 0.146 | 0.688 | 0.081 | |
| Naive Multi-Weak | 0.689 | 0.181 | 0.673 | 0.036 | |
| Joint Decoding | 0.713 | 0.393 | 0.706 | 0.330 | |
| Bayesian Mutli-Weak | 0.700 | 0.276 | 0.697 | 0.252 | |
@@
For the text generation task using slot filling, the performance of the student strong model and teacher weak model finetuned on ground-truth labels is presented in Table 3. The strong ceiling performance (i.e. the performance obtained by training the strong model with ground-truth labels) was 0.781 and the highest weak performance was 0.680.
First, we report Weak(S)-to-Strong performance by using the target sequence generated by the weak model without confidence scores. The results are reported in Table 4. For a single weak model, the Weak-to-Strong performance didn’t necessarily surpass the original weak performance (e.g. OPT-1.3B). Among the models in which Weak-to-Strong performance exceeded the weak performance, the highest PGR was less than 0.15. As for WeakS-to-Strong, the proposed joint decoding approach outperformed both Naive multi-weak and Bayesian multi-weak, achieving the best performance across the table. It can be because both naive multi-weak and Bayesian multi-weak methods use a weak label from each weak model, which potentially includes bad responses. In this case, either using naive multi-weak, which directly trains a model, or using Bayesian multi-weak which trains a model to predict distribution, a bad response can be considered as one of the targets. In contrast, joint decoding allows three weak models to collaboratively determine a single weak label, thereby reducing the chance of selecting a bad response. However, using the auxiliary loss does not provide further improvements in any settings compared to those without due to the lack of confidence measures.
| w/o aux loss | w/ aux loss | ||||
| SLU-F1 | PGR | SLU-F1 | PGR | ||
| Single Weak-to-Strong | GPT2-Large | 0.688 | 0.230 | 0.668 | 0.069 |
| OPT-1.3B | 0.593 | -0.621 | 0.703 | 0.322 | |
| Pythia-1.4B | 0.702 | 0.215 | 0.703 | 0.233 | |
| Naive Multi-Weak | 0.711 | 0.375 | 0.700 | 0.276 | |
| Joint Decoding | 0.706 | 0.330 | 0.706 | 0.330 | |
| Bayesian Mutli-Weak | 0.715 | 0.413 | 0.726 | 0.509 | |
| cDPO | - | - | 0.728 | 0.526 | |
To further improve the performance of text generation, we obtained word-level confidence scores from weak labels for the generated texts and then transformed them into token-level confidence scores that can be used as soft labels to train the student strong model as described in Section 4.2. The results are shown in Table 5. For Weak-to-Strong with a single weak model, all strong models achieved higher PGRs using the proposed soft labels compared to those only with hard labels.
When WeakS-to-Strong is used, consistent performance improvements for all settings were obtained compared to Weak-to-Strong, with the Bayesian multi-weak method achieving the best performance with a PGR of 0.413. The reason why the Bayesian multi-weak method outperformed joint decoding when soft labels were used can be attributed to the fact that the confidence scores help the strong model to learn the confidence of weak models, hence allowing it to differentiate between bad and good sequences. Besides, including the proposed soft labels for WeakS-to-Strong also yielded larger improvements when using both Naive multi-weak and Bayesian multi-weak settings than using hard labels in Table 4. Moreover, with soft labels, WeakS-to-Strong further benefited from using the auxiliary loss, where the strong model performance was increased to 0.726 with a PGR of 0.509.
Based on the strong model supervised finetuned by the Bayesian Multi-Weak approach with auxiliary loss, a cDPO training is performed. The best-performing WeakS-To-Strong scheme is obtained by using Bayesian multi-weak with the auxiliary loss for supervised fine-tuning, followed by training with cDPO, yielding an SLU-F1 of 0.728 with a PGR of 0.526.
@@ 7 Conclusion
This paper extends Weak-to-Strong to WeakS-to-Strong by exploring an ensemble of weak models to simulate the variability in human opinions. A Bayesian inference method, Bayesian WeakS-to-Strong, is proposed to estimate the weak label distribution better using the outputs of existing weak models. Furthermore, the original Weak-to-Strong method can only be applied to text classification tasks, and this paper proposes to extend it to text generation tasks, which allows not only to judge whether content is trustworthy but also to generate trustworthy content. At last, DPO is applied to advance the student model’s preference for learning, beyond the original of teacher-forcing-based learning approach. The results showed the effectiveness of our proposed Bayesian WeakS-to-Strong for both classification and generation tasks, revealing the potential for superalignment.
@@ Appendix A Limitations
The proposed Bayesian WeakS-to-Strong method was tested on two different types of tasks: text classification and generative slot filling. We believe the proposed method is general, and further experiments on other applications are reserved for future work. Due to computational resource limitations, three weak models were used in this paper, mimicking the situation where human annotations are costly and time-consuming to obtain. We believe that the capabilities of the strong model can be recovered to a greater extent when more weak models are involved.
Appendix B Regularising term of EDL
When a sample is not correctly classified, it is expected the total evidence shrinks to zero for the sample. Taking this into consideration, [24] added a regularization term to penalise the misleading evidence. The loss with this regularising term reads
where denotes a Dirichlet distribution with zero total evidence, is the Dirichlet parameter after removal of the non-misleading evidence from predicted , and is the annealing coefficient. By adding a KL-divergence between the Dirichlet distribution with misleading evidence and zero total evidence, the total evidence is enforced to shrink to zero for the simple which is not correctly classified. The annealing coefficient increases by training step, enabling the model to explore the parameter space.
Appendix C Implementation Details
For generation tasks, the AdamW optimizer was used with a linear learning rate scheduler, also with no warm-up. The initial learning rates were set at for GPT2-Large and Pythia-1.4B, and for OPT-1.3B, with a batch size of 8 (mini-batch size of 4). These models were trained for 15 epochs. The checkpoints with the lowest validation loss were selected to ensure the quality of weak labels produced by the weak models. The strong model was trained with a batch size of 2 (mini-batch size of 1) and an initial learning rate of , evaluated at the end of two epochs.
For DPO, the initial learning rate was set to for two epochs, with the cDPO’s hyperparameter set to 10.0 and label smoothing as 0.1. Other settings remain the same as the generation tasks.
Appendix D Broader Impact
@@
Appendix E Licenses for Existing Assets

### v2

@@ 5.2 Models and baselines
For slot filling, Llama-2-7B (Touvron et al., 2023)888https://huggingface.co/meta-llama/Llama-2-7b-hf was used as the strong model which yielded better performance than Qwen-7B in this task. The same set of weak models was used as the classification task. As before, both weak and strong models are finetuned with all model parameters.
Experiments with different numbers of weak models were conducted. One set involved three weak models (GPT2-Large, OPT-1.3B and Pythia-1.4B) in a WeakS-to-Strong experiment, which is referred to as WeakS-to-Strong-3. Another set included all five weak models, called WeakS-to-Strong-5.
Baselines. The proposed Bayesian WeakS-To-Strong method is compared to the following two baselines.
Naive Multi-Weak. The Naive Multi-Weak approach introduced in Section 3.2 servers as the baseline for both classification and generation tasks. For the classification task, the loss can be computed following Eqn. (2). For generation task, Eqn. (2) is modified as follows:
where is the sequence generated by th weak model with length . In contrast to the classification task, where the strong model takes only as input, in the generation task, is paired with the sequence generated by each weak model (which is the weak target) and then fed into the strong model to obtain predictions. Teacher-forcing is used during training.
Joint Decoding. Joint Decoding is used as an additional baseline for text generation tasks. In contrast to the Naive Multi-Weak scheme where each weak model provides a weak target sequence, Joint Decoding employs multiple weak models to collaboratively determine one single target, as illustrated in Figure 1(b). Specifically, we perform Joint Decoding in a re-ranking fashion. For each weak model, the top sequences are generated by beam search in decoding. The sequences from the weak models are gathered to form a list of sequences. Then each sequence is scored in the same way in Section 4.2. The sequence with the highest final score is used as the target sequence for strong model training. Unless otherwise mentioned, was used in the experiments. Each weak-to-strong experiment was run three times with different random seeds. The average and standard deviation of the results were reported. More implementation details can be found in AppendixC.
5.3 Evaluation Metrics
@@
The classification task is evaluated by accuracy, and the SLU-F1 (Bastianelli et al., 2020) is used for slot filling, which combines both word-level and character-level F1 scores to give partial credit to non-exact match predictions. Performance gap recovered (PGR) (Burns et al., 2023) is used to measure the performance gap recovered with weak supervision, which is defined as follows:
@@ 6.1 Text classification
| Weak-to-Strong | GPT2-Large | 0.808 0.007 | 0.503 0.034 | 0.828 0.006 | 0.614 0.029 |
| OPT-1.3B | 0.807 0.005 | 0.541 0.026 | 0.841 0.012 | 0.714 0.058 | |
| Pythia-1.4B | 0.775 0.009 | 0.421 0.042 | 0.793 0.009 | 0.507 0.042 | |
| BLOOM-1B1 | 0.823 0.015 | 0.556 0.087 | 0.843 0.008 | 0.677 0.049 | |
| TinyLlama_v1.1 | 0.832 0.004 | 0.603 0.024 | 0.838 0.005 | 0.643 0.030 | |
| WeakS-to-Strong-3 | Naive Multi-Weak | 0.816 0.002 | 0.586 0.008 | 0.831 0.013 | 0.661 0.064 |
| Bayesian Mutli-Weak | 0.819 0.006 | 0.600 0.033 | 0.850 0.006 | 0.756 0.028 | |
| WeakS-to-Strong-5 | Naive Multi-Weak | 0.832 0.005 | 0.641 0.025 | 0.853 0.006 | 0.754 0.032 |
| Bayesian Mutli-Weak | 0.831 0.008 | 0.627 0.027 | 0.866 0.006 | 0.828 0.038 | |
The proposed Bayesian WeakS-to-Strong approach was first evaluated on a classification task. Table 1 shows the respective performance of the strong model and the weak models trained using ground-truth labels, with the former being the ceiling of the Weak(S)-to-Strong approaches. The strong model has about 7 times the number of parameters as the weak models, which also leads to about 28% relative improvement in the classification accuracy. Results of Weak(S)-to-Strong approaches are shown in Table 2. in Eqn. (1) and Eqn. (5) were set to 0 if auxiliary loss was not used. It can be seen that auxiliary loss was effective for both. Comparing single Weak-to-Strong results, different weak models show significant differences in performance. For example, Pythia-1.4B recovered only 50% of the strong performance, while OPT-1.3B recovered around 70%.
The Naive Multi-Weak approach and Bayesian Multi-Weak using EDL were applied to the classification task for WeakS-to-Strong. Experiments with three weak models (GPT2-Large, OPT-1.3B and Pythia-1.4B) and all five weak models were conducted. With three weak models, the Naive Multi-Weak method got an average PGR of 0.661, which is lower than the single OPT-1.3B model, and Bayesian Multi-Weak boosted PGR to 0.756. This indicates that a naive ensemble approach doesn’t necessarily outperform the best single model, especially when there is a certain weak model which does not perform well. The Bayesian approach can increase the fault tolerance with the usage of prior in this case, as it learns patterns from the entire dataset. With five weak models compared to three, the Bayesian Multi-Weak approach further increased average PGR to 0.828, which is 11% higher than the best single model and 7% higher than the naive baseline, and consistently better with all seeds. The results show the effectiveness of the Bayesian approach for distribution estimation.
@@ 6.2 Text generation
| Strong Model (ceiling) | Llama-2-7B | 6.7B | 0.748 |
@@
| Weak-to-Strong | GPT2-Large | 0.687 0.011 | 0.303 0.125 | 0.673 0.012 | 0.150 0.139 |
| OPT-1.3B | 0.660 0.059 | -0.066 0.715 | 0.696 0.009 | 0.367 0.103 | |
| Pythia-1.4B | 0.702 0.007 | 0.320 0.095 | 0.691 0.022 | 0.173 0.322 | |
| BLOOM-1B1 | 0.690 0.021 | 0.399 0.220 | 0.684 0.021 | 0.337 0.220 | |
| TinyLlama_v1.1 | 0.667 0.006 | 0.000 0.084 | 0.658 0.016 | -0.250 0.224 | |
| WeakS-to-Strong-3 | Naive Multi-Weak | 0.711 0.008 | 0.531 0.101 | 0.694 0.013 | 0.318 0.169 |
| Joint Decoding | 0.703 0.003 | 0.434 0.032 | 0.704 0.015 | 0.442 0.191 | |
| Bayesian Mutli-Weak | 0.712 0.014 | 0.549 0.180 | 0.714 0.014 | 0.574 0.176 | |
| WeakS-to-Strong-5 | Naive Multi-Weak | 0.716 0.010 | 0.606 0.123 | 0.694 0.012 | 0.328 0.144 |
| Joint Decoding | 0.671 0.010 | 0.037 0.120 | 0.675 0.005 | 0.091 0.066 | |
| Bayesian Mutli-Weak | 0.718 0.014 | 0.627 0.173 | 0.721 0.013 | 0.668 0.166 | |
@@
For the text generation task using slot filling, the performance of the student strong model and teacher weak model finetuned on ground-truth labels is presented in Table 3. The strong ceiling performance, which is obtained by training the strong model with ground-truth labels, is 0.748, and the highest weak performance is 0.680.
The Weak-to-Strong and WeakS-to-Strong performances are reported in Table 4. For a single weak model, the Weak-to-Strong model performance didn’t necessarily surpass the original weak performance (e.g. TinyLlama_v1.1), and the highest PGR is 0.399. With five weak models, our proposed Bayesian Multi-Weak approach achieved an average PGR of 0.668, which is 26% better than a single weak model, 6% better than the naive baseline, and consistently better across three seeds. Comparing WeakS-to-Strong performance with and without auxiliary loss, adding auxiliary loss is not effective for naive approaches, but improves the performance of Bayesian approaches. It indicates that the proposed Bayesian method, where the predictions of a strong model are applied as part of the distribution estimation along with weak label as in Eqn. (5), more effectively integrates strong model predictions into the training process, showing the effectiveness of our Bayesian approach for the model ensemble. The ablation study can be found in Appendix D. A noticeable decline has been observed for Joint Decoding when the number of weak models increases from three to five. This may result from the poor performance of the TinyLlama weak model which affects the quality of the generated weak target. Recall that in Joint Decoding, multiple weak models collaboratively determine one single target. In contrast, Naive Multi-Weak and Bayesian Multi-Weak methods still benefit from the increased number of weak models, which indicates that they are more robust against the quality of a single weak model. Additional experiments on Joint Decoding can be found in Appendix F for more analysis and insights.
| SLU-F1 | PGR | |
| Before cDPO | 0.721 0.013 | 0.668 0.166 |
| After cDPO | 0.724 0.015 | 0.705 0.187 |
Based on the strong model supervised by five weak models on the Bayesian Multi-Weak approach with auxiliary loss, a cDPO training is conducted, which further train the model in student-forcing form. Three separate cDPO experiments were conducted using different initial models obtained in SFT stage with three different seeds. The results are shown in Table 5. After cDPO, three models showed consistent performance improvements. The average PGR reached 0.705, 6% relatively better than that before DPO, with a maximum PGR of 0.813.
6.3 Complementarity of weak models
An experiment about the complementarity of different weak models was conducted. For classification models, the agreement is assessed by calculating the accuracy of each model’s predictions on test set, treating outputs from other models as references. For generation models, the Levenshtein distance is calculated between different outputs from two models for a same input, which is obtained using the minimum edit number required to change one sequence to another. The average Levenshtein distance across all samples in the test set is used to measure the agreement between two models.
The results are shown in Figure 3. The agreement among different weak models is among 0.75 for classification models, while for generation task the maximum agreement between different models is 0.56. This suggests that the consistency among different weak models is low, thus they can complement each other well as the faults made by different weak models are not the same. Moreover, comparing with Naive Multi-Weak approach, our proposed Bayesian method estimates a distribution based on weak labels, learning patterns from the entire dataset, thereby increasing the tolerance for fault in weak models, as shown in results in Section 6.1 and 6.2.
@@ 7 Conclusion
This paper extends the Weak-to-Strong framework to WeakS-to-Strong by leveraging an ensemble of weak models to capture the variability in human opinions. We propose a Bayesian inference method, Bayesian WeakS-to-Strong, to more accurately estimate the weak label distribution based on the outputs of multiple weak models. Additionally, while the original Weak-to-Strong method was limited to text classification tasks, this paper expands its applicability to text generation, enabling both the assessment of content trustworthiness and the generation of trustworthy content. Finally, DPO is utilized to enhance the student model’s preference for learning, going beyond the traditional teacher-forcing approach. Our results demonstrate the effectiveness of Bayesian WeakS-to-Strong for both classification and generation tasks, highlighting its potential for superalignment.
@@ Appendix A Limitations
The proposed Bayesian WeakS-to-Strong method was tested on two different types of tasks: text classification and generative slot filling. We believe the proposed method is general, and further experiments on other applications are reserved for future work. Due to computational resource limitations, experiments on only three and five weak models were conducted in this paper, mimicking the situation where human annotations are costly and time-consuming to obtain. We believe that the capabilities of the strong model can be recovered to a greater extent when more weak models are involved.
Appendix B Broader Impact
@@
Appendix C Implementation Details
For generation tasks, the AdamW optimizer was used with a linear learning rate scheduler, also with no warm-up. The initial learning rates were set at for GPT2-Large and Pythia-1.4B, and for OPT-1.3B, with a batch size of 8 (mini-batch size of 4). These models were trained for 15 epochs. The checkpoints with the lowest validation loss were selected to ensure the quality of weak labels produced by the weak models. The strong model was trained with a batch size of 2 (mini-batch size of 1) and an initial learning rate of , evaluated at the end of two epochs. In the training for strong ceiling performance, the hyperparameters were adjusted based on the validation set. For Weak-to-Strong training, in which case the ground truth is not accessible, we aligned the hyperparameter settings with those used in strong ceiling training.
For DPO, the initial learning rate was set to for two epochs, with the cDPO’s hyperparameter set to 2.0 and label smoothing as 0.1. Other settings remain the same as the generation tasks.
Appendix D Ablation Study on Probability Estimation for Weak Sequences
The importance of probability estimation for weak sequences was explored in this section. Two key steps were calculating target wordpiece probability using the word as a gap bridge, and estimating probabilities of other categories by scaling the strong output (see Section 4.1 for details). The results are shown in Table 6. The average PGR of using a one-hot label (Line 1) is negative, showing that WeakS-to-Strong performance doesn’t surpass the weak model without probability estimation. Estimating probability for the target wordpiece significantly improved performance compared to using one-hot labels. This indicates the necessity of introducing target probabilities during training, which allows the strong student model to learn the weak model’s confidence for its generated labels, considering it may be incorrect. Based on this, adding probabilities of other categories, the overall precision of target label estimation improves, resulting in about 20% PGR gain. This experiment highlights the importance of precise probability estimation for weak sequences.
| Target wordpiece | Other category | SLU-F1 | PGR |
| 0.656 0.016 | -0.145 0.204 | ||
| ✓ | 0.704 0.025 | 0.453 0.305 | |
| ✓ | ✓ | 0.721 0.013 | 0.668 0.166 |
Appendix E Weights of Weak Models
The selection for weights of weak models was explored in this section, where experiments on average, dynamic and fixed weights were conducted. For average, the weight in Eqn. (2) and Eqn. (4) is set to the same value for each weak model. That is to say, the losses are calculated for each weak model respectively and then averaged as the final loss. For fixed weight, the weights are set to a fixed value based on the performance of different weak models. For dynamic weight, for a certain sample, the confidence in the utterance of a weak model is treated as a weight for that weak model, considering the importance of different weak models could vary on different samples. The results are shown in Table 7. Compared to the average, fixed weight improves the average PGR by about 8% because it assigns weights based on the performance of different weak models, rather than just averaging. Dynamic weight shows a noticeable decline in performance compared to fixed weight. It may be because our proposed per-token target probability already provided the student model with fine-grained information about the reliability of the label, making dynamic weight entirely redundant.
| SLU-F1 | PGR | |
| Average | 0.715 0.019 | 0.589 0.237 |
| Dynamic weight | 0.687 0.013 | 0.241 0.163 |
| Fixed weight | 0.721 0.013 | 0.668 0.166 |
Appendix F Performance of Joint Decoding
F.1 The Impact of the Quality of Weak Models
This section provides additional experiments where different weak model combinations are used: (i) three weak models (GPT2-Large, OPT-1.3B, Pythia-1.4B), (ii) four models (adding BLOOM-1B1); and (iii) all five models (further adding TinyLlama_v1.1). For TinyLlama_v1.1, two setups were investigated: (i) only using it to generate weak target sequence but not for scoring (scoring done by other four models); (ii) using it for both generation and scoring. The results are listed in Table 8. Compared to the three-weak-model system, introducing the th model shows a 10% increase in PGR. However, incorporating TinyLlama as the th weak model undermines the performance (last row in Table 8), even nearly fails to surpass weak model performance. A possible reason is that TinyLlama performs poorly on the text generation task. The impact of the TinyLlama can be reduced by excluding it from scoring. As shown in the second to the last row of Table 8, without TinyLlama scoring, the results with five weak models are close to those with three. This indicates that TinyLlama’s poor scoring ability prevents it from selecting a good target sequence among the candidates and the poor quality from a certain weak model can largely impact the overall results for Joint Decoding methods. In contrast, as discussed in Section 6.2, the proposed Bayesian WeakS-to-Strong method is more robust against the quality of a single weak model.
F.2 The Impact of Beam Size
As introduced in Section 5.2, in Joint Decoding, each weak model generates output sequences by beam search in decoding. This section investigates Joint Decoding with different beam sizes , as shown in Table 9. For WeakS-to-Strong-3, results with different yields similar results. However, for WeakS-to-Strong-5, with TinyLlama included in the weak model set, the result with is significantly worse than that with and . The experiments on beam size further demonstrate that the poor results of the WeakS-to-Strong-5 are due to its weak ability to select the target sequence, as introduces more distractions than .
| 3 Weak models | BLOOM-1B1 | TinyLlama_v1.1 | SLU-F1 | PGR | |
| Generating | Scoring | ||||
| ✓ | 0.703 0.003 | 0.434 0.032 | |||
| ✓ | ✓ | 0.711 0.008 | 0.549 0.100 | ||
| ✓ | ✓ | ✓ | 0.706 0.009 | 0.452 0.086 | |
| ✓ | ✓ | ✓ | ✓ | 0.671 0.010 | 0.037 0.120 |
| WeakS-to-Strong-3 | WeakS-to-Strong-5 | |||
| SLU-F1 | PGR | SLU-F1 | PGR | |
| M=3 | 0.706 0.012 | 0.471 0.155 | 0.669 0.023 | 0.022 0.281 |
| M=5 | 0.703 0.003 | 0.434 0.032 | 0.671 0.010 | 0.037 0.120 |
| M=10 | 0.703 0.004 | 0.431 0.045 | 0.650 0.004 | -0.224 0.050 |
Appendix G Regularising Term of EDL
When a sample is not correctly classified, it is expected the total evidence shrinks to zero for the sample. Taking this into consideration, Sensoy et al. (2018) added a regularization term to penalise the misleading evidence. The loss with this regularising term reads
where the KL term refers to the in Section 3.3, denotes a Dirichlet distribution with zero total evidence, is the Dirichlet parameter after removal of the non-misleading evidence from predicted , and is the annealing coefficient. By adding a KL-divergence between the Dirichlet distribution with misleading evidence and zero total evidence, the total evidence is enforced to shrink to zero for the simple which is not correctly classified. The annealing coefficient increases by training step, enabling the model to explore the parameter space.
Appendix H Prompt Used in Experiment
Appendix I Licenses for Existing Assets